show that (fog)(x)=(gof)(x)=x
f(x)=4x; g(x)=1/4 x

Answer by jim_thompson5910(35256)

f(x) = 4x

f(g(x)) = 4(g(x))

f(g(x)) = 4((1/4x)

f(g(x)) = (4*1/4)x

f(g(x)) = (4/4)x

f(g(x)) = 1*x

f(g(x)) = x

(f o g)(x) = x
--------------------------------------------

g(x) = (1/4)x

g(f(x)) = (1/4)(f(x))

g(f(x)) = (1/4)(4x)

g(f(x)) = ((1/4)*4)(x)

g(f(x)) = (4/4)x

g(f(x)) = 1*x

g(f(x)) = x

(g o f)(x) = x